OpenStudy (anonymous):

Solve the problem. A company makes three chocolate candies: cherry, almond, and raisin. Matrix A gives the amount of ingredients in one batch. Matrix B gives the costs of ingredients from suppliers X and Y. What is the cost of 100 batches from supplier X? pictures down below in a moment

OpenStudy (anonymous):

A=

OpenStudy (anonymous):

B=3 2 sugar 3 4 choc 2 2 milk

OpenStudy (anonymous):

A. $6600 B. $7800 C. $3300 D. $4800

OpenStudy (mandre):

This looks like matrix multiplication. \[\left[\begin{matrix}4 & 6 & 1 \\ 5&3&1\\ 3 & 3 & 1\end{matrix}\right].\left[\begin{matrix}3 &2 \\ 3&4\\ 2 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

im so confused by it all...

OpenStudy (anonymous):

please help lol

OpenStudy (mandre):

It's 3x3 times 3x2 so you'll end up with a 3x2 a11 (That's the number in row 1, column 1) = 4*3+6*3+1*2 (row * column). a21(row 2, column 1) = 5*3+3*3+1*2 You with me so far?

OpenStudy (anonymous):

uh no.. :/

OpenStudy (mandre):

You don't know matrix multiplication?

OpenStudy (anonymous):

no...

OpenStudy (mandre):

Then you have a problem. I presume this is Linear algebra?

OpenStudy (anonymous):

yea im on a exam atm and im stuck on this question but i really need the points for it tho

OpenStudy (mandre):

Sorry, then I can't help you any further.